OpenStudy (anonymous):

What is the debt to total assets ratio? Current assets $16,000 Noncurrent assets 22,875 Total assets $38,875 Current liabilities $11,349 Long-term debt 10,599 Total stockholders' equity 16,927 Total liabilities and stockholders' equity $38,875 Cash from operating activities $5,468 Cash paid for capital expenditures 4,160 Dividends paid 538 Average shares outstanding 869

OpenStudy (gw2011):

(11,349+10,599)/38,875=21,948/38,875=0.5646=56.46%